• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 10:20:30 on March 4,2005

    Entry number 140831

    Followups:

    keyword=DAQ restoration

    In collab. with others, the DAQ was restored:

    1. Cables to R-HRS plugged in. Someone should check coinc. time.
    2. Power supplies on L-arm VME and floor TS fixed.
    3. S2m-Left cable plugged into scalers. I think this was out for
    a long time (weeks ?).

    The experiment should run cosmics and do a pulser check of
    coincidences. Page me this weekend if problems.